### v1

@@ Abstract
In robot-assisted minimally invasive surgery, we introduce the Surgical Action Planning (SAP) task, which generates future action plans from visual inputs to address the absence of intraoperative predictive planning in current intelligent applications. SAP shows great potential for enhancing intraoperative guidance and automating procedures. However, it faces challenges such as understanding instrument-action relationships and tracking surgical progress. Large Language Models (LLMs) show promise in understanding surgical video content but remain underexplored for predictive decision-making in SAP, as they focus mainly on retrospective analysis. Challenges like data privacy, computational demands, and modality-specific constraints further highlight significant research gaps. To tackle these challenges, we introduce LLM-SAP, a Large Language Models-based Surgical Action Planning framework that predicts future actions and generates text responses by interpreting natural language prompts of surgical goals. The text responses potentially support surgical education, intraoperative decision-making, procedure documentation, and skill analysis. LLM-SAP integrates two novel modules: the Near-History Focus Memory Module (NHF-MM) for modeling historical states and the prompts factory for action planning. We evaluate LLM-SAP on our constructed CholecT50-SAP dataset using models like Qwen2.5 and Qwen2-VL, demonstrating its effectiveness in next-action prediction. Pre-trained LLMs are tested zero-shot, and supervised fine-tuning (SFT) with LoRA is implemented to address data privacy concerns. Our experiments show that Qwen2.5-72B-SFT surpasses Qwen2.5-72B with a 19.3% higher accuracy. The source code and dataset will be available at https://anonymous.4open.science/r/SAP
@@ 1 Introduction
In Robot-assisted Minimally Invasive Surgery (RMIS), current intelligent applications, such as surgical workflow recognition [2, 11, 5], instrument segmentation [1, 20], action recognition [13, 4], and medical question-answering [3, 16], primarily focus on retrospective analysis rather than future decision-making. While valuable for procedure analysis, they lack the ability to support intraoperative future decision-making, underscoring the need for our Surgical Action Planning (SAP) task, a forward-looking framework. Our SAP streamlines complex procedures by decomposing them into discrete actions, and generates long-horizon sequential action plans from visual inputs, enabling the achievement of user-defined objectives. For instance, natural language directives like “Conduct Calot’s triangle dissection, duct and vessel clipping, and dissection from the liver bed" can be converted into a sequence of executable surgical actions, providing procedural guidance and demonstrating the potential for procedure automation.
@@ 2 Methods
2.1 Surgical Action Planning Task (SAP) Definition
@@ 2.2.2 Prompts Factory for Generating Action Plans
Fig. 1) shows the prompts factory, including (1) Descriptive captions prompts (DCPrompts) To generate descriptive captions using the VLM , we employ the following prompt: “You are a professional surgical analysis assistant specializing in laparoscopic cholecystectomy. Your task is to generate detailed descriptions of the video frame provided, focusing on anatomical structures, tool manipulation, key surgical steps, and environmental features." (2) Action planning prompts (APPrompts): Firstly, Surgical Domain Knowledge Base is build based on surgical process, safety protocol, and action description. Next, the APPrompts include “Based on the provided Surgical Domain Knowledge Base, Descriptions of near frames and the current state (IndirNHFM) OR Video frames provided (DirNHFM), Previous actions, Last action, you need to provide the following: Progress Assessment, Safety Considerations, Ready-to-Execute Actions: Provide three actions. For each action, provide a rationale explaining why it is ranked in that order."
@@ 3.0.2 Metrics
We report the sample-level accuracy and video-level accuracy under both standard and relaxed conditions.
@@ 3.0.3 Action planning results
The analysis of the experimental results can be summarized as follows: (1) IndirNHFM vs. DirNHFM: In zero-shot experiments, Qwen2.5-72B achieves a standard top-1 SLAcc of 45.61 and VLAcc of 53.42 with IndirNHFM, outperforming Qwen2-VL with DirNHFM by 5.26% (45.61 vs. 40.35) in SLAcc and 5.05% (53.42 vs. 48.37) in VLAcc. This indicates that indirect visual observation better captures contextual nuances, potentially due to richer textual representations. (2) Zero-Shot vs. SFT: Qwen2.5-72B-SFT outperforms Qwen2.5-72B, achieving 19.3% higher standard top-2 SLAcc (78.95 vs. 59.65), 20.74% higher standard top-2 VLAcc (80.05 vs. 59.31), 9.3% higher relaxed top-2 SLAcc (93.02 vs. 83.72), and 4.12% higher relaxed top-2 VLAcc (93.33 vs. 89.21). This demonstrates that SFT significantly boosts model performance for SAP by leveraging task-specific data. (3) Qwen2.5-32B-SFT (best model) vs others: Although Qwen2-VL-72B-SFT achieves the best standard top-1 SLAcc of 47.37 and VLAcc of 54.31, Qwen2.5-32B-SFT demonstrates strong overall performance across all metrics, with 44.3% higher standard top-2 SLAcc (78.95 vs. 54.39), 41.8% higher standard top-2 VLAcc (83.05 vs. 58.55), 28.1% higher relaxed top-2 SLAcc (95.35 vs. 74.42), and 16.2% higher relaxed top-2 VLAcc (96.67 vs. 83.25).

### v2

@@ Abstract
In robot-assisted minimally invasive surgery, we introduce the Surgical Action Planning (SAP) task, which generates future action plans from visual inputs to address the absence of intraoperative predictive planning in current intelligent applications. SAP shows great potential for enhancing intraoperative guidance and automating procedures. However, it faces challenges such as understanding instrument-action relationships and tracking surgical progress. Large Language Models (LLMs) show promise in understanding surgical video content but remain underexplored for predictive decision-making in SAP, as they focus mainly on retrospective analysis. Challenges like data privacy, computational demands, and modality-specific constraints further highlight significant research gaps. To tackle these challenges, we introduce LLM-SAP, a Large Language Models-based Surgical Action Planning framework that predicts future actions and generates text responses by interpreting natural language prompts of surgical goals. The text responses potentially support surgical education, intraoperative decision-making, procedure documentation, and skill analysis. LLM-SAP integrates two novel modules: the Near-History Focus Memory Module (NHF-MM) for modeling historical states and the prompts factory for action planning. We evaluate LLM-SAP on our constructed CholecT50-SAP dataset using models like Qwen2.5 and Qwen2-VL, demonstrating its effectiveness in next-action prediction. Pre-trained LLMs are tested in a zero-shot setting, and supervised fine-tuning (SFT) with LoRA is implemented. Our experiments show that Qwen2.5-72B-SFT surpasses Qwen2.5-72B with a 19.3% higher accuracy.
@@ 1 Introduction
In Robot-assisted Minimally Invasive Surgery (RMIS), current intelligent applications, such as surgical workflow recognition [2, 11, 5], instrument segmentation [1, 20], action recognition [13, 4], and medical question-answering [3, 16], primarily focus on retrospective analysis rather than future decision-making. While valuable for procedure analysis, they lack the ability to support intraoperative future decision-making, underscoring the need for our Surgical Action Planning (SAP) task, a forward-looking framework. Our SAP streamlines complex procedures by decomposing them into discrete actions, and generates long-horizon sequential action plans from visual inputs, enabling the achievement of user-defined objectives. For instance, natural language directives like “Conduct Calot’s triangle dissection, clipping and cutting of duct and vessel, and dissection of the gallbladder from the liver bed" can be converted into a sequence of executable surgical actions, providing procedural guidance and demonstrating the potential for procedure automation.
@@ 2 Methods
2.1 Surgical action planning (SAP) task definition
@@ 2.2.2 Prompts factory for generating action plans
Fig. 1) shows the prompts factory, including (1) Descriptive captions prompts (DCPrompts) To generate descriptive captions using the VLM , we employ the following prompt: “You are a professional surgical analysis assistant specializing in laparoscopic cholecystectomy. Your task is to generate detailed descriptions of the video frame provided, focusing on anatomical structures, tool manipulation, key surgical steps, and environmental features.” (2) Action planning prompts (APPrompts): Firstly, surgical domain knowledge base is build based on surgical process, safety protocol, and action description. Next, the APPrompts include “Based on the provided Surgical Domain Knowledge Base, Descriptions of near frames and the current state (IndirNHFM) OR Video frames provided (DirNHFM), Previous actions, Last action, you need to provide the following: progress assessment, safety considerations, ready-to-execute actions: Provide three actions. For each action, provide a rationale explaining why it is ranked in that order.”
@@ 3.0.2 Metrics
We report the sample and video-level accuracy under standard and relaxed conditions.
@@ 3.0.3 Action planning results
The analysis of the experimental results can be summarized as follows: (1) IndirNHFM vs. DirNHFM: In zero-shot experiments, Qwen2.5-72B achieves a standard top-1 SLAcc of 45.61% and VLAcc of 53.42% with IndirNHFM, outperforming Qwen2-VL with DirNHFM by 5.26% (45.61% vs. 40.35%) in SLAcc and 5.05% (53.42% vs. 48.37%) in VLAcc. This indicates that indirect visual observation better captures contextual nuances, potentially due to richer textual representations. (2) Zero-Shot vs. SFT: Qwen2.5-72B-SFT outperforms Qwen2.5-72B, achieving 19.3% higher standard top-2 SLAcc (78.95% vs. 59.65%), 20.74% higher standard top-2 VLAcc (80.05% vs. 59.31%), 9.3% higher relaxed top-2 SLAcc (93.02% vs. 83.72%), and 4.12% higher relaxed top-2 VLAcc (93.33% vs. 89.21%). This demonstrates that SFT significantly boosts model performance for SAP by leveraging task-specific data. (3) Qwen2.5-32B-SFT (best model) vs others: Although Qwen2-VL-72B-SFT achieves the best standard top-1 SLAcc of 47.37 and VLAcc of 54.31, Qwen2.5-32B-SFT demonstrates strong overall performance across all metrics, with 44.3% higher standard top-2 SLAcc (78.95% vs. 54.39%), 41.8% higher standard top-2 VLAcc (83.05% vs. 58.55%), 28.1% higher relaxed top-2 SLAcc (95.35% vs. 74.42%), and 16.2% higher relaxed top-2 VLAcc (96.67% vs. 83.25%).